# Building spatial world models from sparse transitional episodic memories

### v1

@@ 4.7 ESWM Enables Fast Adaptation to Changes
Unlike traditional world models that embed environment knowledge in their weights, ESWM decouples memory from reasoning. This separation makes ESWM inherently adaptable to changes: when the environment changes, no retraining is needed—only the external memory bank is updated. Because ESWM operates on sparse, one-step transitions, targeted edits to memory are straightforward. We demonstrate this flexibility in a dynamic navigation task. After introducing substantial changes—such as new obstacles—an ESWM agent continues to navigate between arbitrary states with a 93% success rate (Fig.5c).
The adaptive agent works as follows: Given a (possibly outdated) memory bank $M$, it plans a path from $s_{\text{start}}$ to $s_{\text{goal}}$ (Section  4.6). During execution, it compares predicted observations with actual ones. When mismatches occur (indicating environmental changes), the agent removes outdated transitions (where $s_{s}$ or $s_{e}$ match the predicted observation from $M$), prompting exploration (Section  4.5) to fill the gaps. After acquiring new transitions, the agent replans with the updated memory.
Compared to a baseline LSTM agent trained via RL to navigate between any two states in a fixed environment (Extended Method A.7) and EPN (see Section D), ESWM is far more robust to change. For example, after adding a new wall consisting of multiple contiguous obstacles, ESWM maintains a 93% navigation success rate, whereas EPN and the baseline drop to 72% and 56% respectively (see Fig.5c). Unsurprisingly, the baseline agent fails to form any structured latent spatial map.
@@ 5 Discussion
In this work, we introduced ESWM, a neural network model that can rapidly construct a coherent spatial world model from sparse and fragmented one-step transitions. ESWM exhibits strong spatial reasoning and supports downstream tasks like exploration and navigation—even without explicit training for them. Unlike prior models that rely on fixed circuitry for spatial representation (52; 5; 33), ESWM learns both representation and update mechanisms directly from experience. It encodes new memories in a single shot, avoiding the iterative parameter tuning required by earlier approaches (52; 5; 47). By storing transitional memories instead of state-observation pairs (54; 9; 55), ESWM also adapts quickly to environmental changes, such as the addition or removal of obstacles—an area where many existing models struggle.
@@
While our focus was on a grid environment where episodic memories are transitions between integer observations via allocentric actions (e.g, ’go north’), preliminary results on MiniGrid (6) (Fig.A9) show that ESWM’s prediction performance and internal map representation generalize to a more complex setting with pixel observations and egocentric actions (e.g, ’turn left’). Yet still, such formulation of episodic memories greatly simplifies the complex sensory observations and egocentric motion cues typically associated with humans’ — a gap we leave for future work. Another limitation is the use of highly controllable and simplified environments, which allowed us to scale up the training and control different aspects of the environment and data freely. Future work could explore extending this approach to more realistic settings with natural, complex sensory information. Although theoretical, our work has potential societal impacts in autonomous navigation and exploration, highlighting the need for safe, ethical, and responsible deployment (see Section E).
Appendix A Extended Methods
A.1 Training Details
Transistion Embedding.
@@ A.1 Training Details
In the Open Arena environment, each state value is encoded as a 6-bit binary. Each bit is independently projected—via either the source-state or end-state linear projector—into a 128-dimensional embedding. The six embeddings are concatenated into a single 768-dimensional state vector. Actions are similarly mapped into 768-dimensional embeddings using a dedicated linear projector. The final transition embedding is then obtained by averaging the three 768-dimensional vectors corresponding to (1) the start-state, (2) the action, and (3) the end-state.
@@
In the Random Wall environment, the same procedure applies, except that each state is represented by an integer and projected into a 1,024-dimensional embedding, yielding a 1,024-dimensional transition vector.
Model Architectures.
@@
  • Transformer: 8 attention heads per layer; feed-forward hidden dimension of 2,048.
@@
Training Hyperparameters.
@@
A.2 Building Random Environments
@@
  • $A$: Maps an undirected edge $e=(l_{i},l_{j})\in E$, which is a pair of locations, to an action-induced transition either starting from $l_{i}$ or $l_{j}$ with equal chances. e.g. $A(l_{i},l_{j})$ returns a 3-tuple transition ($l_{i}$, go West, $l_{j}$) or ($l_{j}$, go East, $l_{i}$) with equal probability.
@@
A.3 Generating Memory Banks
@@
    Use $A$ to map edges in $T$ to directed 3-tuple transitions between locations and $\psi$ to map locations to observations. This produces an array of transitions between observations $[(s_{\text{s}},a,s_{\text{e}})]$.
@@ A.4 Query Selection
The model receives the $M$ and $q=(s_{\text{s}},a,s_{\text{e}})$, with either $s_{\text{s}}$, $a$, or $s_{\text{e}}$ masked randomly with equal probability, as input, and asked to infer the masked compartment.
A.5 ISOMAP
We consider ESWM’s population neuron activity at a particular location as the ESWM’s activation vector as it infers the missing component of a query transition that either starts or ends at that location. For end-state and action predictions (i.e, source state is visible to ESWM), the activation vector is associated with the source state’s location, while for start-state prediction (i.e, source state is not visible), we associate the activation with the end state’s location. For transformer models, we used the query token after each transformer encoder block as the activation vector, while for multilayer LSTM and Mamba, we used the final hidden state (i.e, the last token from the processed sequence) from the last layer.
To collect ESWM’s activation across all locations in an environment, we task ESWM to do either start-state, action, or end-state prediction for all transitions in the environment conditioned on a memory bank. This process is repeated with multiple memory banks from the same environment until approximately 1,000 activation vectors are collected. We then apply ISOMAP using cosine distance and 20–60 neighbors to generate the visualizations.
A.6 Euclidean Distance Estimation From Activation
@@
A.7 Baseline Navigational Agent
The baseline agent is trained via A2C [Sutton, 2018, Mnih, 2016] to navigate between any two states in a single Random Wall environment, which differs from ESWM and EPN, which are meta-trained across many environments. At each time step $t$, the agent receives as input a sequence of historical interactions with the environment, each as a tuple ($s_{t-1},a_{t-1},s_{t},s_{\text{goal}}$), and is tasked to output an action towards the goal state $s_{\text{goal}}$. Each part of the tuple is distinctly embedded onto a 128D vector and the vectors are averaged to produce the final observation at each time step. The embedded sequence is fed into a one-layer LSTM (with a hidden dimension of 256) and the predicted action and value are read out from the final hidden state via distinct linear heads.

### v2

@@ 4.6 ESWM Enables Fast Adaptation to Changes
Unlike traditional world models that embed environment knowledge in their weights, ESWM decouples memory from reasoning. This separation makes ESWM inherently adaptable to changes: when the environment changes, no retraining is needed—only the external memory bank is updated. Because ESWM operates on sparse, independent one-step transitions, targeted edits to memory are straightforward. We demonstrate this flexibility in a dynamic navigation task. After introducing substantial changes—such as new obstacles—an ESWM agent continues to navigate between arbitrary states with a 93% success rate (Fig. 5c).
The adaptive agent works as follows: Given a (possibly outdated) memory bank $M$, it plans a path from $s_{\text{start}}$ to $s_{\text{goal}}$ (Section  4.5). During execution, it compares predicted observations with actual ones. When mismatches occur (indicating environmental changes), the agent removes outdated transitions (where $s_{s}$ or $s_{e}$ match the predicted observation from $M$), prompting exploration (Section  4.4) to fill the gaps. After acquiring new transitions, the agent replans with the updated memory.
ESWM’s adaptability significantly surpasses both EPN and a baseline RL agent trained specifically for the original environment (see Extended Methods C.10 for baseline details). For example, after adding a new wall consisting of multiple contiguous obstacles, ESWM maintains a 93% navigation success rate, whereas EPN and the baseline drop to 72% and 56% respectively (see Fig. 5c). Unsurprisingly, the baseline agent fails to form any structured latent spatial map.
4.7 Scaling ESWM
In the previous sections, we focused on simple grid environments, where episodic memories are transitions between integer observations via allocentric actions (e.g., go north), to examine how ESWM behaves under various scenarios. To test ESWM’s scalability, we next evaluate it in Minigrid and ProcThor (see Section3.3)–two more complex scenarios with egocentric high-dimensional observations and actions. To adapt ESWM to MiniGrid and ProcThor, its operation on a discrete bank of transitions remains unchanged, and we only adjusted the encoding function for states (trained from scratch for MiniGrid, pretrained vision model for ProcThor) and actions and the readout function for generating the predictions to match those in this environment (Extended Methods C.1).
In MiniGrid, ESWM continues to exhibit strong predictive performance (Fig. 6a) Top-Right) while constructing coherent internal map representations (Fig. 6a) Bottom) that adapt to the environment’s structure. In ProcThor, our results show that ESWM is solving this challenging task by learning to predict both state vectors and action parameters (Fig. 6)b).
@@ 5 Discussion
In this work, we introduced ESWM, a NN model that can rapidly construct a coherent spatial world model from sparse and fragmented one-step transitions. ESWM exhibits strong spatial reasoning and supports downstream tasks like exploration and navigation—even without explicit training for them. Unlike prior models that rely on fixed circuitry for spatial representation (Wang et al., 2023; Chandra et al., 2025; Kymn et al., 2024), ESWM learns both representation and update mechanisms directly from experience. It encodes new memories in a single shot, avoiding the iterative parameter tuning required by earlier approaches (Wang et al., 2023; Chandra et al., 2025; Stachenfeld et al., 2017). By storing transitional memories instead of state-observation pairs (Whittington et al., 2020; Coda-Forno et al., 2022b; Whittington et al., 2022), ESWM also adapts quickly to environmental changes, such as the addition or removal of obstacles—an area where many existing models struggle.
@@
A limitation is the use of highly controllable environments, which allowed us to scale up the training and control different aspects of the environment and data freely. Future work could explore extending this approach to more realistic settings such as robotic navigation with natural, complex sensory information. Although theoretical, our work has potential societal impacts in autonomous navigation and exploration, highlighting the need for safe, ethical, and responsible deployment (see Section I).
6 Reproducibility Statement
We provide detailed descriptions to ensure the reproducibility of our work. These include training details (Extended Methods C.1, data generation pipeline (Extended Methods C.2 - C.4), baseline implementations (Extended Methods C.5, C.10, Section H), methods of analysis (Extended Methods C.6, C.7), and algorithm pseudo codes (Section G).
7 Declaration of The Use of LLM
We recognize the importance of LLM usage disclosure and declare that our use of LLM is strictly limited to word editing and polishing.
8 Acknowledgement
M.D. was supported by the UNIQUE PhD Fellowship. This research was supported by the NSERC Discovery grant RGPIN-2021-03035, and CIHR Project Grant PJT-191957. P.B. was supported by FRQ-S Research Scholars Junior 1 grant 310924, FRQNT-NSERC NOVA grant 2024-NOVA-346823, and the William Dawson Scholar award. All analyses were executed using resources provided by the Digital Research Alliance of Canada (Compute Canada) and funding from Canada Foundation for Innovation project number 42730.
Appendix
Appendix A Relation to other models of medialxx temporal lobe
Successor Representations (SR). In the context of reinforcement learning, successor representation has been shown to explain many non-spatial aspects of hippocampus responses (Stachenfeld et al., 2017). SR and ESWM are similar in that both rely on transition information, but they differ in several important ways. First, classical SR methods require either a full transition matrix or incremental experience sufficient to estimate one. They are not designed to infer missing transitions or to complete a partially observed graph. ESWM, by contrast, is built specifically for that setting: it takes a sparse set of one-step episodic transitions and infers the missing structural relations directly at inference time. Conceptually, ESWM is estimating the entire family of transition distributions from incomplete data, whereas SR presupposes access to the full underlying dynamics.
Second, SRs are learned through iterative, policy-dependent updates and encode multi-step, discounted future occupancy. ESWM does not rely on iterative updates and is not tied to a particular policy. The transitions in its memory bank are single, policy-free episodic observations, and the model performs transitive reasoning over these episodes on demand to answer arbitrary transition queries.
In addition, SRs collapse the environment into a future-occupancy representation, which mixes structural and predictive information. ESWM keeps the structure explicit and manipulable, allowing it to reconstruct the transition graph itself rather than a discounted predictive summary (e.g. state values). ESWM is also symmetric in its inference capabilities: it jointly estimates forward transitions, backward transitions, and actions from the same memory bank, whereas SRs typically encode only forward predictive relationships.
Finally, ESWM supports rapid structural updates by directly editing or adding episodic transitions, without needing to relearn a transition matrix. This contrasts with SRs and related models, which must re-estimate transition dynamics when the environment changes.
Tolman-Eichenbaum Machine. Tolman–Eichenbaum Machine (TEM)(Whittington et al., 2020; 2022; Bakermans et al., 2025) has been proposed as a model of medial temporal lobe function and is meta-trained across diverse environments to predict unseen transitions from episodic experience. TEM factorizes sensory and structural information: the shared structural scaffold is encoded in the weights of a recurrent network, while environment-specific observations are bound to that scaffold through a Hebbian-slot memory or self-attention module. The recurrent state therefore encodes a generic structural prior—typically a Euclidean, grid-like topology—and the attractor dynamics allow for content-specific retrieval. This makes TEM well-suited for environments that share this underlying family of structures and allows it to tolerate large changes in sensory input at familiar locations. However, this same design makes the model less flexible when the underlying topology itself changes or when the environment obeys different structural rules (e.g., RandomWall environments where walls can appear in arbitrary shapes and configurations).
In contrast, ESWM does not factorize structural and sensory information. Both the structural layout and the expected sensory observations are inferred jointly and directly from episodic transitions. Because ESWM does not rely on a fixed structural template encoded in its weights, it can accommodate changes to the environment by directly modifying its memory bank–allowing it to update predictions immediately without retraining. This makes ESWM particularly effective in settings where the topology can vary widely across environments or can change abruptly within an environment.
Finally, whereas TEM operates on sequential trajectories and depends on an explicit dynamical update of its recurrent state, ESWM integrates disjoint and non-sequential experiences. This enables more sample-efficient learning from sparse observations and supports rapid adaptation to newly introduced obstacles or structural changes that were never part of a trajectory.
Recurrent Predictive Models. Recurrent predictive models have also been proposed as models of neural computations within the hippocampus (Levenstein et al., 2024). These models consist of recurrent neural networks trained to predict future observations from a sequence of prior observations along a particular trajectory. In these models, both the structural and sensory information are embedded directly in the recurrent network’s weights. This architecture makes them even less flexible in the face of changes, more so than the TEM class of models. Whereas TEM includes an episodic module that allows for rapid adaptation to sensory changes within a fixed structural template, recurrent predictive models must update their recurrent weights to incorporate both sensory and structural changes, which is considerably harder to achieve quickly.
In addition, recurrent predictive models depend on having continuous, sequential experience to gradually shape the internal dynamics into a coherent predictive structure. This reliance on dense experience makes them poorly suited to integrating disjoint or fragmentary observations. ESWM, by contrast, is designed specifically to operate on sparse, non-sequential episodic transitions, allowing it to infer the structural relationships among states without needing long trajectories.
Appendix B Evidence for building spatial maps from disjoint experiences in animals
Several behavioral studies have demonstrated that animals and humans can integrate information collected across disjoint episodes into coherent spatial or relational maps. (Roberts et al., 2007) showed that rats inferred a shortcut between two locations in a maze after exploring its subcomponents in separate sessions on different days. From the perspective of our framework, the animals observed disconnected edges of the underlying spatial graph—never experiencing them as a continuous trajectory—yet were still able to infer the globally shortest route during the test. (Fernandez et al., 2023) similarly demonstrated that humans trained to navigate locally within three separate environments on different days could integrate these fragmented experiences to navigate across them globally. (Park et al., 2020) found that human subjects who learned each dimension of a 2D relational space in separate episodes nonetheless formed a joint map of the full space, again illustrating cross-episode integration.
There is also substantial neural evidence that hippocampal and amygdala circuits integrate memories according to shared structure rather than temporal contiguity. (Yokose et al., 2017) showed that amygdala-dependent memories from distinct tasks form overlapping engrams when they share a critical stimulus. (Schlichting et al., 2015) found that the anterior hippocampus integrates across paired associates (A→B, B→C) to support the inferred A→C relationship. (McKenzie et al., 2014) reported that hippocampal population codes become similar for objects that share spatial context, even when those objects were experienced in separate episodes. Related reviews (Morton et al., 2017; Brunec et al., 2020) synthesize a broad literature showing that the hippocampus organizes memories into integrated representations that reflect spatial, temporal, and conceptual structure—often beyond what was directly and continuously experienced.
Altogether, these findings provide strong evidence that biological systems do face and solve a form of the ”disjoint memory” problem. They routinely combine episodic fragments acquired at different times, in different contexts, and along different task dimensions, to construct coherent internal maps that support flexible behavior.
Appendix C Extended Methods
C.1 Training Details
C.1.1 Open Arena
@@
In the Open Arena environment, each integer state value is encoded as a 6‑bit binary. Each bit is independently projected—via either the source‑state or end‑state linear projector—into a 128‑dimensional embedding. The six embeddings are concatenated into a single 768‑dimensional state vector. Actions are similarly mapped into 768‑dimensional embeddings using a dedicated linear projector. The final transition embedding is then obtained by averaging the three 768‑dimensional vectors corresponding to (1) the start‑state, (2) the action, and (3) the end‑state.
C.1.2 Random Wall
@@
In the Random Wall environment, the same procedure applies, except that each state is represented by an integer and projected into a 1,024‑dimensional embedding, yielding a 1,024‑dimensional transition vector.
C.1.3 MiniGrid
In the Minigrid environment, the states are 5 $\times$ 5 egocentric view, observing 25 locations all at once. Each location in the environment is assigned one of the colored objects (e.g, red floor) among a pre-defined set of 9 possible colored objects. Each of the objects among the 25 observed objects in a state is encoded into a 64 dimensional vector separately. The encodings of the 25 objects are then concatenated to produce a single 1600 dimensional representation for the state. Similar to Open Arena and Random Wall, the $s_{s}$, $a$, and $s_{e}$ embeddings are averaged to produce the final embedding for the transition. The prediction for the state is 25 independent 9-way classifications.
C.1.4 ProcThor Experiments
In ProcThor (Deitke et al., 2022), an agent can be placed anywhere in the environment to collect $224\times 224\times 3$ RGB observations (i.e. states; FOV=90). A transition consists of a start state $s_{s}$, a continuous action vector $\Delta x,\Delta y,\Delta\theta]$ specifying how much to move right, forward, and rotate clockwise (negative values mean moving along the opposite axis) to arrive at the end state, and an end state $s_{e}$. To sample a transition, we first spawn the agent at a random, unobstructed location with a random rotation (start state). Then we randomly teleport the agent to a new location that is within the $3\times 3$m box of the start state, followed by a random rotation (end state). Lastly, we compute the egocentric action vector from the start state to the end state’s poses.
The states are embedded by DINOV3 (Siméoni et al., 2025) into $W\times H\times D=14\times 14\times 1024$ feature map, padded and pooled into $W\times H\times D=4\times 4\times 1024$, then flattened into 16, 1024-dimensional tokens. The action is projected by a 1-layer MLP into a standalone 1024-dimensional token. The start state, end state, and action tokens are concatenated into a sequence of 33 tokens to represent a single transition. Suppose a memory bank has $N$ memories, then ESWM receives the concatenation of the memory bank and query, totaling $(N+1)\times 33$ tokens, as input.
The action and end state of the query transition are masked with equal chances (the start state is left out for our preliminary experiment). For state prediction, the model is tasked to maximize the cosine similarity between its output and DINOV3’s representation. For action prediction, the model performs categorical prediction over a discretized action space. To discretize action prediction, the $3m\times 3m$ movement box is discretized into grids, with each grid representing a class. ESWM needs to essentially predict how many steps to move right/left, forward/backward until the end state’s location is reached, with the step size (or discretization granularity) determined so that there are 64 classes. The rotation is discretized by 10-degree intervals, resulting in 36 classes.
The concatenation of the memory bank and query is fed into a multi-layer transformer. After all transformer blocks, we linearly read out the action prediction from the action token, and directly match the state tokens with their target representations without any further transformation. We apply cross-entropy loss to action predictions and cosine distance loss to states.
ProcThor Training Pipeline
To train ESWM in GPU-based environments like ProcThor, we devise a memory-efficient, fast, and scalable training pipeline that largely follows DD-PPO (Wijmans et al., 2020). The training is distributed over $N$ GPUs, with each GPU alternating between online data generation across $K$ scenes and training. In our preliminary experiment, we train ESWM with 4 A100 GPUs for 19 days over 80k iterations with an effective batch size of 144, using only 250 GB of memory.
C.1.5 Model Architectures.
@@ C.1.4 ProcThor Experiments
  • Transformer: 8 attention heads per layer; feed‑forward hidden dimension of 2,048, unless otherwise specified.
@@
For ProcThor experiment, we replaced the standard transformer block with the Alternating Attention block from VGGT (Wang et al., 2025). Each transformer block consists of a memory-wise attention layer, where the attention scope is limited to the tokens within a transition – consolidating the relation between the start and end state, and a global attention layer, where all tokens can attend to each other – enabling memory integration. We use 16 heads in our ProcThor runs.
C.1.6 Training Hyperparameters.
@@
C.2 Building Random Environments
@@
  • AA: Maps an undirected edge $e=(l_{i},l_{j})\in E$, which is a pair of locations, to an action-induced transition either starting from $l_{i}$ or $l_{j}$ with equal chances. e.g. $A(l_{i},l_{j})$ returns a 3-tuple transition ($l_{i}$, go West, $l_{j}$) or ($l_{j}$, go East, $l_{i}$) with equal probability.
@@
C.3 Generating Memory Banks
@@
    Use $A$ to map edges in $T$ to directed 3-tuple transitions between locations and $\psi$ to map locations to observations. This produces an array of transitions between observations $(s_{\text{s}},a,s_{\text{e}})]$.
@@ C.4 Query Selection
The model receives the $M$ and $q=(s_{\text{s}},a,s_{\text{e}})$, with either $s_{\text{s}}$, $a$, or $s_{\text{e}}$ masked randomly with equal probability, as input, and asked to infer the masked component.
C.5 TEM-T Implementation
Our goal is to maximally equalize the training of TEM-T and ESWM to eliminate potential confounds such as task differences. Thus, TEM-T also receive the concatenation of memory bank and unseen query transition as input, except that the disjoint memory bank is replaced with a sequential trajectory that spans the entire environment. The task is to predict the missing end-state of the final query transition, analogous to inferring the masked end state in ESWM. The trajectory is kept minimal, generated by NetworkX traveling salesman problem, to match the sparsity of the memories we provide for ESWM. In Random Wall, we explicitly ensure that the agent runs into each obstacle at least once throughout the traversal so that the model has sufficient knowledge to infer the underlying structure of the environment.
To our best knowledge, no publicly available code exists for TEM-T. Therefore, we resort to a custom implementation that closely follows the authors’ recipe. We use the same dimensions to encode the positions $e_{t}$ and observations $x_{t}$ (384 in Open Arena and 512 in Random Wall) and ensure the combined dimension of $h_{t}=[x_{t},e_{t}]$ matches ESWM’s embedding dimension. Sigmoid is used as the non-linear activation $\sigma$ in positional encoding generation, following the observation that linear and ReLU activation produce exploding gradients. We use 8 heads and 2048 feed-forward dimensions, matching the configuration of ESWM. The prediction is read out linearly from the embedded query transition after a single/multiple transformer blocks. We train TEM-T over 460k iterations with a batch size of 128 and a cosine learning rate schedule, same as ESWM.
C.6 ISOMAP
We consider ESWM’s population neuron activity at a particular location as the ESWM’s activation vector as it infers the missing component of a query transition that either starts or ends at that location. For $s_{e}$ and $a$ predictions (i.e, $s_{s}$ is visible to ESWM), the activation vector is associated with the $s_{s}$’s location, while for $s_{s}$ prediction (i.e, $s_{s}$ is not visible), we associate the activation with the $s_{e}$’s location. For transformer models, we used the query token after each transformer encoder block as the activation vector, while for multilayer LSTM and Mamba, we used the final hidden state (i.e, the last token from the processed sequence) from the last layer.
To collect ESWM’s activation across all locations in an environment, we task ESWM to do either $s_{s}$ $a$, or $s_{e}$ prediction for all transitions in the environment conditioned on a memory bank. This process is repeated with multiple memory banks from the same environment until approximately 1,000 activation vectors are collected. We then apply ISOMAP using cosine distance and 20–60 neighbors to generate the visualizations.
C.7 Euclidean Distance Estimation From Activation
@@
C.8 sRSA
We use the formula from (Levenstein et al., 2024) to compute sRSA. This metric measures the Spearman correlation between latent and physical distances. To get the latent representation at a particular location, we task the model to solve all queries that start in that location and average the activation. We repeat this procedure across all locations in an environment and across 5 memory banks from each environment. We used ESWM-T-2L’s first layer’s query token’s activation vector during the action prediction task for this analysis.
C.9 Noisy Memory Banks Implemention
To create memory banks with noisy, or contradictory, state observations, we pick a location and change the corresponding memories in the memory bank to observe different states at that location.
To create memory banks with noisy, or contradictory, transitions, we randomly pick a transition from the memory bank, edit its end state, and append the modified transition to the memory bank.
This analysis is done with ESWM-T-4L in Random Wall.
C.10 Baseline Navigational Agent
The baseline agent is trained via A2C (Sutton, 2018; Mnih, 2016) to navigate between any two states in a single Random Wall environment, which differs from ESWM and EPN, which are meta-trained across many environments. At each time step $t$, the agent receives as input a sequence of historical interactions with the environment, each as a tuple ($s_{t-1},a_{t-1},s_{t},s_{\text{goal}}$), and is tasked to output an action towards the goal state $s_{\text{goal}}$. Each part of the tuple is distinctly embedded onto a 128D vector and the vectors are averaged to produce the final observation at each time step. The embedded sequence is fed into a one-layer LSTM (with a hidden dimension of 256) and the predicted action and value are read out from the final hidden state via distinct linear heads.